There are 30 students in the group. The marks for the control test were the following: 8 students received “excellent” marks, 12 students - “good” marks, 8 students – “satisfactory”. What is the probability that 3 students invited for answer have “good” marks


1
Expert's answer
2021-09-12T23:51:58-0400
"P(3\ ''good'')=\dfrac{\dbinom{12}{3}\dbinom{30-12}{0}}{\dbinom{30}{3}}"

"=\dfrac{\dfrac{12!}{3!(12-3)!}(1)}{\dfrac{30!}{3!(30-27)!}}=\dfrac{220}{4060}=\dfrac{11}{203}"